Exercise 08-3 Preparing flexible budgets LO P1 Tempo Companys fixed budget (based on sales of 12,000 units) for the first quComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Compute the toComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Compute the toComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Compute the inComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Compute the in
Answer #1
Variable Amount
per Unit
Total Fixed Cost
Sales 204
Variable costs:
Direct materials 23 =276000/12000
Direct labor 43 =516000/12000
Production supplies 27 =324000/12000
Sales commissions 7 =84000/12000
Packaging 15 =180000/12000
Total variable costs 115
Contribution margin 89
Fixed costs:
Plant manager salary 76,000
Advertising 100,000
Administrative salaries 126,000
Depreciation—office equip. 96,000
Insurance 66,000
Office rent 76,000
Total fixed costs 540,000
1
Variable cost per unit 115
2
Total fixed costs 540000
3
Income from Operations at sales of 10,000 units 350000 =(10000*89)-540000
4
Income from Operations at sales of 14,000 units 706000 =(14000*89)-540000